OpenStudy (anonymous):

in rolling two dice, what is the probability that the sum of the dice is 9 given that the sum of the dice is at least 3?

OpenStudy (anonymous):

i get the answer 0.1111

OpenStudy (kropot72):

The probability that one of the dice will have a three is 1/6 and on the same throw the probability the other will have a 6 is also 1/6. Therefore the two events are independent the probability of the sum being 9 is the product of the two probabilities.\[P(3\cup 9)=\frac{1}{6}\times \frac{1}{6}=\frac{1}{36}\] The same applies for the other 3 combinations that result in a sum of 9.

OpenStudy (kropot72):

The sum of the four probabilities of 1\36 = 4\36 = 0.1111

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)